### v3

@@ 8.3 Forecasting Uncertainty
As a statistical matter, high classification accuracy can support high forecasting accuracy. However, in addition to the complications just described, we picked the true heat wave dates before the classification analysis began. Our results are conditional on first knowing the dates of the true heat wave and the faux heatwaves and then assigning the appropriate outcome classes homogeneously to all grid cells. That alone makes forecasting problematic. In particular, one cannot properly take the next step into valid forecasting.
One should ordinarily expect proper forecasts coupled with estimates of uncertainty. A good option for discrete outcomes is conformal prediction sets that produce legitimate forecasts and valid uncertainty estimates even in finite samples. The key assumption is exchangeability (Gupta et al., 2022). Under the popular split samples approach, the data are randomly separated into two disjoint samples: (1) a training dataset that a classifier can use for fitting, and (2) a calibration dataset used construct prediction sets. A wide variety of classifiers can be used be they algorithms or models. There is no requirement that either contains all of the relevant predictors or employs the substantively appropriate functional forms. Still, for unlabeled cases needing a forecast, a conformal prediction set will include a forecast at a predetermined coverage probability properly interpreted as the probability that the forecast is correct. No assumptions need be made about the joint probability distribution of the dataset or additional properties of the calibration data.
But, one must be able to make a credible case that at least the calibration data are exchangeable. This is difficult requirement for the PNW data. Because we selected the outcomes in advance, the realized data are not IID or exchangeable, and the calibration data would not be exchangeable either. In addition, the algorithm specification incorporated by the classifier is the product of various kinds of prior analyses raising post-model-selection inferential challenges (Berk et al., 2013; Kuchibhotla et al., 2022). But it is still useful to consider whether the results of the exploratory data analysis are replicable.
@@ 9 A Replication Study in Phoenix, Arizona
Repeating a forecasting exercise in a new geographical setting can have several benefits. First, one is able to consider whether the statistical approaches employed earlier generalize well. Does the formulation of a particular data generating process more broadly apply? Are gain scores instructive elsewhere? Are the random forests and the GA genetic algorithm appropriate in other locations? Second, there is substantive generalizability. In particular, using the same variables with their values realized elsewhere, how well reproduced is the confusion table from the Pacific Northwest? All of these questions can be properly addressed with no feature engineering, no “bake-offs” between algorithms, no model selection, no p-hacking, and no filtering of reported results.
@@ 9.1 Data Challenges
We obtained AIRS daily data for the Phoenix area for 63 grid cells from 2019 through 2023. Data were not available for 2024. The measured variables were the very same used in the PNW analysis. However, the climate and topography in the greater Phoenix area is very different from the PNW and far more homogeneous. The Pacific Northwest is lush. The greater Phoenix area is largely a desert. We picked the greater Phoenix area precisely because of the strong replication challenges it posed.
We had about half the number of grid cells for the Phoenix area than for the Pacific Northwest. There were also a substantial number of grid cells with missing data. Between the two, we anticipated a much smaller number of observations overall. One result could much weaker statistical power coupled with still more sparsity. We needed a procedure to increase the number of usable grid cells for the within-subject (i.e., grid cells) approach.
@@
All three time periods were then stacked to make single data set. Each grid cell in principle was observed in three different years and in principle, there would be 189 observations overall, each exposed to one set of extreme heat wave precursors and two sets of faux heat wave precursors. One benefit is that the collection of study subjects was the same under all three conditions; the could be no changes in sample composition.
@@
But the potential problems from missing data remained. When we included only the three precursor variables required for comparability that were used for the PNW forecasts, the number of grid cells with complete data fell to 128. Because the data were lost through cloud cover and occasional downtime on the remote sensing satellite, one might argue that the data were missing in blocks conditionally at random. Nevertheless, the large amount of missing data raises concerns about the representativeness of the Phoenix data and about substantial instabilities that could be introduced. There would be further complications should there be differential amounts of non-random missingness when the grid cells were exposed to true heat wave precursors versus the when the grid cells were exposed to the faux heat wave precursors. The design structure is shown in Figure 7.171717 Because the data were likely generated in an unknown, spatially dependent matter, all of the missing data imputation methods with which were familiar seem inappropriate.
@@ 9.2 Comparisons Between True Heat Wave and Faux Heat wave Predictor Histograms
As before, we compare the distributions for each of three predictors when exposed to the true heat wave precursors compared to the faux heat wave precursors. The amount of missing data is far more pronounced for the true heat wave data. We expected about 33% of the data to be from the faux heat wave cases, and about 67% of the cases to be from the faux heat wave cases. After listwise deletion, 27% of the observations were from the true heat wave, and 73% were from the faux heat waves. The numbers of observations were 35 and 93 respectively. True heat wave cases are noticeably underrepresented, and the sample size is small enough to be worrisome.
For the true heat wave and faux heat waves, Figure 8 shows the top two plots different predictor distributions for tropopause height and for temperature at level 8. But compared to Phoenix, the true heat wave predictor values tend to be toward the left of the faux heat wave values. The bottom plot for the water vapor mixing ratio, is again inconclusive. Compared to the PNW data, one might speculate that the larger fraction of missing values from the true heat wave precursors are disproportionately from greater tropopause heights and higher temperatures. That might account for some results that at this point seem counterintuitive.
@@ 9.3 Classification Results
We began simply. Using the random forests classifier already trained for with data from the Pacific Northwest, we obtained the fitted probabilities with input data from the greater Phoenix area. That is, the trained algorithm was fixed. The data from Phoenix were employed solely to construct fitted probabilities. Whatever the relationships captured for the PNW were being implicitly assumed to apply in Phoenix.
The resulting Phoenix confusion table has an overall misclassification proportion of a little over .16, which is about 3 times larger the overall misclassification proportion for the PNW. There are no false positives, but the false negative fraction is 60%. This substantial difference in the proportion of false negatives compared to the proportion of false positives may be in part a result of the marginal distribution of the response variable. Whereas there are for the PNW data nearly the same number of actual heat wave precursor exposures as faux heat wave precursor exposures, the proportions for Phoenix have approximately two faux heat wave exposures for each actual heat wave exposure. With weighting methods described in the next section, the empirical ratio of false negatives to false positives could be made to approximate 1 to 1, but the complications that can follow would need to be addressed.
A second comparability assessment was undertaken using the random forest classifier just as before with the Phoenix data structured as in Figure 7. All three three predictors used in the PNW analysis are used again. After deletion of all rows with any missing data, there were 128 observations. With so few observations in each, substantial instability was expected. The implicit assumption was that the same precursors would matter, but probably not in the same way as before.
We were pleasantly surprised by the confusion table produced. The overall misclassification rate for Phoenix is a little over .03, about the same as error rate for the PNW. There were no false positives, and a false negative rate of approximately .11 The difference was likely explained in part by the relative decline the number of in the heat wave cases, now accounting for just a little over a quarter of the data. All three predictors contribute substantially to the fit, although the role of atmospheric temperature 14 days earlier is more dominant than for the PNW.181818 Given the likely instability because of the modest number of training data observations, we examined the confusion tables following from 10 additional random splits of the data. Overall misclassification proportions ranged from a little over .05 to a little under .12. Often there were no false positives. By most standards, classification accuracy was quite good but lacked the stability desired.
@@
Despite the histogram differences just described and great differences in climate and topography, Figure 6 for the PNW and Figure 9 for the greater Phoenix area have their smoothers showing much the same S-shaped relationship between the fitted probabilities and atmospheric temperature at an altitude of 8.191919 As before, the vertical variation in the plots is visually attenuated. Both relationships begin flat for their lowest values before their slopes turns positive, and end flat at the highest values. The Phoenix relationship is shifted a bit to the right in response to its higher temperatures. It also somewhat less smooth because of data that are more sparse.
Comparing the PNW and Phoenix smooths for tropopause height, both start out flat before turning positive. For the PNW, the overall pattern is roughly linear and positive thereafter. For the greater phoenix area, the relationship smoothly turns negative at around 15,500 meters, and the tropopause heights start as somewhat higher values.
@@
The apparent contradiction between the predictor comparative histogram outcomes and the partial dependence plots is perhaps explained by the difference between marginal and conditional distributions. The partial dependence plots are conditioned so that adjustments are made for the dependence between the three predictors. For example, in the Phoenix data, the correlation between temperature at level 8 and tropopause height is .83.
Despite the excellent classification accuracy, one might wonder if there remain dependencies in the Phoenix analysis. Given the spatial nature of the data, one might be curious about spatial dependence between grid cells implying relationships not yet addressed. Classifier residuals were computed. The spatial correlogram R procedure gstat, authored and maintained by Edzer Pebesma, was applied. Figure 10 shows the result. The correlogram shows that the measure of association, here Moran’s I, fluctuates around 0.0 and is not systematically related to the distance between spatial residuals.202020 The residuals can in principle range between ±$\pm 1.0$. The discrete unit for the plot is .01. The horizontal axis is in distance multiples of .01. After a Bonferroni correction for multiple testing, for none of the associations is the null hypothesis of no association rejected. The nominal critical values was .05.
@@
It is important to emphasize that although the PNW analysis was steeped in extensive data exploration, the Phoenix analysis had none. Yet classification accuracy is similar for both, and two of the three partial dependence plots are roughly similar in both locations. Moreover, one might think that for Phoenix, conformal prediction sets can now be used. As before, however, we knew the days on which the Phoenix heat wave occurred before the data analysis began and as before, grid cells in advance were coded homogeneously. When grid cells were exposed to true heat wave precursors, they were all code equal to 1.0. When grid cells were exposed to faux heat wave precursors, they were all code equal to 0.0. The data are neither IID nor exchangeable.
@@ 10 Problems from Endogenous Sampling
The cross-over designs for the PNW and for the greater Phoenix area led to excellent balance that remedied the rare events problem for our binary outcome of heat wave or not. However, it is challenging to think of a naturally occurring population of grid cells that could be a plausible source of the data in which such balance would have occurred. Our curation process substantially oversampled extreme heat wave days, which is a form of troublesome endogenous oversampling.
@@
For numeric response variables, clever sampling strategies have been proposed that, in effect, oversample for extreme values (Ragone et al., 2018). Endogenous sampling is explicit and worrisome. For discrete response variables, our focus here, several related remedies have been proposed (Cosslett, 1993; Waldman, 2000). A dataset can be weighted in different ways so that the imbalance is corrected and the endogenous (response) variable once again can provide a valid estimates. Sometimes the weighting is buried in the statistical machinery along with other performance enhancements, but it is still altering the sample imbalance. For example, when Matthew’s correlation, also know as the phi coefficient, is used as an objective function for a classifier to maximize, it can help correct imbalance (Dourard it al., 2024). But because the cost ratio target of false positives to false negative is set to 1.0, the marginal distribution is, in effect, being re-weighted.
@@
There are several difficulties with such solutions. To begin, in many applications it is not apparent what the proper population is or even should be. An earlier example was whether the population for studying a rare extreme heat wave should be daily ARIS data for the most recent 5 years, daily AIRS data for the most recent year, or daily AIRS data for the most recent month. The choice can depend heavily how heat wave forecasting will be undertaken in practice. If essential responses to extreme heat waves are determined month to month, monthly data might be preferred. But then, a trained forecasting algorithm for each month would be needed, perhaps updated year to year. The work load could be reduced if forecasters were prepared to assume that rare extreme head waves only occur in the summer months. The forecasters would then be backing into the kinds of designs we used in the PNW and greater Phoenix area.
@@
In response to asymmetric costs, there are weighting adjustments leading to what Kuhn and Johnson (2013: Chapter 6) call “cost sensitive training.” Related thinking can be found in focal loss approaches (Lin et al., 2017) and implicitly in gradient boosting (Friedman, 2001). Rednick and his colleagues (2021) offer the most involved approach that weights more heavily the hard to classify observations, and a specified outcome class to can be given a higher priority as well.222222 We coded their loss function in keras3 in R. It ran, but did far worse that the usual cross-entropy loss, perhaps because the classifier was already too successful with the within-subject design and the balanced data that resulted. There were too few false negatives and false positives usefully to weight. We also found tuning a conceptual challenge when we tried to return to a highly imbalanced source population response variable and with a plausible stakeholder consensus cost ratio. In their classic book, Luce and Raiffa (1957) formulate the problem as risk minimization rather than loss minimization. Broadly framed, asymmetric loss functions are required to account for asymmetric fitting costs.
@@
The conflation of corrections for sample imbalance with the estimated misclassification cost ratio is easily illustrated. Using the expressions above for the weights, the PNW forecasting exercise was repeated. Recall that the response variable was binary, and there were the same number of grid cells that experienced a heat wave as did not experience a heat wave. Using the month of June, 2021 as the target population, for a little more than 13% of the days, the 144 grid cells experienced the extreme heat wave. For the other days, the 144 grid cells did not. In the data analyzed, for 50% of the days, the 144 grid cells experienced the extreme heat wave. A disparity between 13% and 50% may be more extreme than anticipated by Manski and Lerman (1977), but we can proceed for illustrative purposes.
@@
After weighting, the population imbalance was approximately reproduced. But that’s not all. Before weighting, the classification error rate was a little over 5%, and there were nearly the same number of false negatives as false positives. The cost ratio was about 1.0, conveying equal costs. When Manski-Lerman weighing was applied, the classification error rate rose to nearly 14%, and the cost ratio of false positives to false negatives increased from about 1.0 to nearly 10.0. False positives were now almost 10 times more costly than false negatives. In other words, the costs of heat wave false alarms were substantially up-weighted compared to the costs of true alarms not given. Many stakeholders might strongly prefer the reverse. In short, weighting changes the prior distribution of the response variable and the relative costs of false negatives to false positives.
@@
For the forecasting results reported for the PNW and the greater Phoenix area, we had no access to stakeholders nor any concrete policy context. The numbers of false negatives and false positives were very few and of roughly equal size. Classifier training determined an empirical cost ratio of about 1 to 1. Going forward in a real policy setting combined with stakeholder input, the within-subject research design could be revised to better capture practical concerns. The main requirement is that grid cells continue to capitalize on the within-subject structure. Additional faux heat waves could easily be added so that a more appropriate, but manageable, imbalance was achieved. Another possibility might be to train the classifier on balanced data but to improve fitted values, use as input an altered dataset with many more faux heat waves.

### v4

@@ 8.3 Forecasting Uncertainty
As a statistical matter, high classification accuracy can support high forecasting accuracy. However, in addition to the complications just described, we picked the true heat wave dates before the classification analysis began. Our results are conditional on first knowing the dates of the true heat wave and the faux heatwaves and then assigning the appropriate outcome classes homogeneously to all grid cells. When the grid cells had been exposed two weeks earlier to extreme heat wave precursors, each grid cells was coded as “1.” When the grid cells had been exposed two weeks earlier to faux heat wave precursors, each grid cells was coded as “0.” That makes forecasting even more problematic. In particular, one cannot properly take the next step into valid forecasting.
One should ordinarily expect proper forecasts coupled with estimates of uncertainty. A good option for discrete outcomes is conformal prediction sets that produce legitimate forecasts and valid uncertainty estimates even in finite samples. The key assumption is exchangeability (Gupta et al., 2022). Under the popular split samples approach, the data are randomly separated into two disjoint samples: (1) a training dataset that a classifier can use for fitting, and (2) a calibration dataset used construct prediction sets. A wide variety of classifiers can be used be they algorithms or models. There is no requirement that either contains all of the relevant predictors or employs the appropriate functional forms. Still, for unlabeled cases needing a forecast, a conformal prediction set will include a forecast at a predetermined coverage probability properly interpreted as the probability that the forecast is correct. No assumptions need be made about the joint probability distribution of the dataset or additional properties of the calibration data.
But, one must be able to make a credible case that at least the calibration data are exchangeable. This is difficult requirement for the PNW data. Because we selected the outcomes in advance, the realized data are not IID or exchangeable, and the calibration data would not be exchangeable either. In addition, the algorithm specification incorporated by the classifier is the product of various kinds of prior analyses raising post-model-selection inferential challenges (Berk et al., 2013; Kuchibhotla et al., 2022).
@@ 9 A Replication Study in Phoenix, Arizona
Repeating the clasification exercise in a new geographical setting can have several benefits. First, one is able to consider whether the statistical approaches employed earlier generalize well. Does the formulation of a particular data generating process more broadly apply? Are gain scores instructive elsewhere? Are the random forests and the GA genetic algorithms appropriate in other locations? Second, there is substantive generalizability. In particular, using the same variables with their values realized elsewhere, how well reproduced is the confusion table from the Pacific Northwest? All of these questions can be properly addressed with no feature engineering, no “bake-offs” between algorithms, no model selection, no p-hacking, and no filtering of reported results.
@@ 9.1 Data Challenges
We obtained AIRS daily data for the Phoenix area for 63 grid cells from 2019 through 2023. Data were not available for 2024. The measured variables were the very same used in the PNW analysis. However, the climate and topography in the greater Phoenix area are very different from the PNW and far more homogeneous. The Pacific Northwest is lush. The greater Phoenix area largely is a desert. We picked the greater Phoenix area because of the strong replication challenges it posed.
We had about half the number of grid cells for the Phoenix area than for the Pacific Northwest. There were also a substantial number of grid cells with missing data. Between the two, we anticipated a much smaller number of observations overall. One result could much weaker statistical power coupled with still more sparsity. We sought a procedure to increase the number of usable grid cells for the within-subject (i.e., grid cells) approach.
@@
All three time periods were then stacked to make single data set. Each grid cell in principle was observed in three different years and in principle, there would be 189 observations overall, each exposed to one set of extreme heat wave precursors and two sets of faux heat wave precursors. One benefit is that the collection of study subjects was the same under all three conditions; there could be no changes in sample composition.
@@
But the potential problems from missing data remained. When we included only the three precursor variables required for PNW comparability, the number of grid cells with complete data fell to 128. Because the data were lost through cloud cover and occasional downtime on the remote sensing satellite, one might argue that the data were missing in blocks, conditionally at random. Nevertheless, the large amount of missing data raises concerns about the representativeness of the Phoenix data and about substantial instabilities that could be introduced. There would be further complications should there be differential amounts of non-random missingness when the grid cells were exposed to true heat wave precursors versus the when the grid cells were exposed to the faux heat wave precursors. The design structure is shown in Figure 7.171717 Because the data were likely generated in an unknown, spatially dependent matter, all of the missing data imputation methods with which were familiar seem inappropriate.
@@ 9.2 Comparisons Between True Heat Wave and Faux Heat wave Predictor Histograms
As before, we compare the distributions for each of three predictors when exposed to the true heat wave precursors compared to the faux heat wave precursors. The amount of missing observations are more pronounced for the true heat wave data. We expected about 33% of the data to be for the true heat wave precursors, and about 67% of the data to be for the faux heat wave precursors. After listwise deletion, the proportions were 27% and 73% respectively. The corresponding numbers of observations were 35 and 93. True heat wave cases are noticeably underrepresented, and the sample size is small enough to be worrisome.
For the true heat wave and faux heat waves, Figure 8 shows in the top two plots different predictor distributions for tropopause height and for temperature at level 8. But in contrast to the PNW, the true heat wave predictor values tend to fall toward the left of the faux heat wave values. The bottom plot for the water vapor mixing ratio, is again inconclusive. Compared to the PNW data, perhaps because of the larger fraction of missing values from the true heat wave precursors, greater tropopause heights and higher temperatures are disproportionately lost from the red bars. That might account for some results that at this point seem counterintuitive.
@@ 9.3 Classification Results
We began simply. Using the random forests classifier already trained with data from the Pacific Northwest, we obtained the fitted probabilities with input data from the greater Phoenix area. That is, the trained algorithm was fixed. The data from Phoenix were employed solely to construct fitted probabilities. Whatever the relationships captured for the PNW were being implicitly assumed to apply in Phoenix.
The resulting Phoenix confusion table has an overall misclassification proportion of a little over .16, which is about 3 times larger the overall misclassification proportion for the PNW. There are no false positives, but the false negative fraction is 40%. This substantial difference in the proportion of false negatives compared to the proportion of false positives may be in part a result of the marginal distribution of the response variable. Whereas there are for the PNW data nearly the same number of actual heat wave precursor exposures as faux heat wave precursor exposures, the proportions for Phoenix have approximately two faux heat wave precuresor exposures for each actual heat wave precursor exposure. With weighting methods described in the next section, the empirical ratio of false negatives to false positives could be made to approximate 1 to 1, but the complications that can follow would need to be addressed.
A second comparability assessment was undertaken using the random forest classifier just as before, but with the Phoenix data structured as in Figure 7. All three predictors used in the PNW analysis are used again. After deletion of all rows with any missing data, there were 128 grid cell observations. With so few observations, substantial instability was expected. The implicit assumption was that the same precursors would matter, but probably not in the same way as before.
We were pleasantly surprised by the confusion table produced. The overall misclassification rate for Phoenix is a little over .03, about the same as error rate for the PNW. There were no false positives, and a false negative rate of approximately .11 The difference was likely explained in part by the relative decline in the number of in the heat wave cases, now accounting for just a little over a quarter of the data. All three predictors contribute substantially to the fit, although the role of atmospheric temperature 14 days earlier is more dominant than for the PNW.181818 Given the likely instability because of the modest number of training data observations, we examined the confusion tables following from 10 additional random splits of the data. Overall misclassification proportions ranged from a little over .05 to a little under .12. Often there were no false positives. By most standards, classification accuracy was quite good but lacked the stability desired.
@@
Despite the histogram differences just described and great differences in climate and topography, the partial dependence plots in Figure 6 for the PNW and for Figure 9 for the greater Phoenix area have smoothers showing much the same S-shaped relationship between the fitted probabilities and atmospheric temperature at an altitude of 8.191919 As before, the vertical variation in the plots is visually attenuated. Both relationships begin flat for their lowest values before their slopes turns positive, and end flat at the highest values. The Phoenix relationship is shifted a bit to the right in response to its higher temperatures. It also somewhat less smooth because of data that are more sparse.
Comparing the PNW and Phoenix smooths for tropopause height, both start out flat before turning positive. For the PNW, the overall pattern is roughly linear and positive thereafter. For the greater Phoenix area, the relationship smoother turns negative at around 15,500 meters, and the tropopause heights start as somewhat higher values.
@@
The apparent contradiction between the predictor comparative histograms and the partial dependence plots is perhaps explained by the difference between marginal and conditional distributions. The partial dependence plots are conditioned so that adjustments are made for the dependence between the three predictors. For example, in the Phoenix data, the correlation between temperature at level 8 and tropopause height is .83.
Despite the excellent classification accuracy, one might wonder if there remain dependencies in the Phoenix analysis. Given the spatial nature of the data, there could remain spatial dependence between grid cells implying relationships not yet addressed. Classifier residuals were computed. The spatial correlogram R procedure gstat, authored and maintained by Edzer Pebesma, was applied. Figure 10 shows the result. The correlogram shows that the measure of association, here Moran’s I, fluctuates around 0.0 and is not systematically related to the distance between spatial residuals.202020 The residuals can in principle range between ±$\pm 1.0$. The discrete unit for the plot is .01. The horizontal axis is in distance multiples of .01. After a Bonferroni correction for multiple testing, for none of the associations is the null hypothesis of no association rejected. The nominal critical values was .05.
@@
It is important to emphasize that although the PNW analysis was steeped in extensive data exploration, the Phoenix analysis had none. Yet classification accuracy is similar for both, and two of the three partial dependence plots are roughly similar in across the two locations. But as before, conformal prediction sets cannot properly be used. Again, we knew the days on which the Phoenix heat wave occurred prior to the data analysis, and grid cells in advance were coded homogeneously. The data are neither IID nor exchangeable.
@@ 10 Problems from Endogenous Sampling
The within-grid-cell designs for the PNW and for the greater Phoenix area led to excellent balance that remedied the rare events problem for our binary outcome of heat wave or not. However, it is challenging to think of a naturally occurring population of grid cells that could be a plausible source of the data in which such balance would have occurred. Our curation process substantially oversampled extreme heat wave days, which is a form of endogenous oversampling.
@@
For numeric response variables, clever sampling strategies have been proposed that, in effect, oversample for extreme values (Ragone et al., 2018). Endogenous sampling is explicit and worrisome. For discrete response variables, our focus here, several related remedies have been proposed (Cosslett, 1993; Waldman, 2000). A dataset can be weighted in different ways so that the imbalance is corrected, and the endogenous (response) variable once again can provide a valid estimates. Sometimes the weighting is buried in the statistical machinery along with other performance enhancements, but it is still altering the sample imbalance. For example, when Matthew’s correlation, also know as the phi coefficient, is used as an objective function for a classifier to maximize, it can help correct imbalance (Dourard it al., 2024). But because the cost ratio target of false positives to false negative is set at 1.0, the marginal distribution is, in effect, being re-weighted.
@@
There are several difficulties with such solutions. To begin, in many applications it is not apparent what the proper population is or even should be. An earlier example was whether the population for studying a rare extreme heat wave should be daily ARIS data for the most recent 5 years, daily AIRS data for the most recent year, or daily AIRS data for the most recent month.
@@
In response to asymmetric costs, there are weighting adjustments leading to what Kuhn and Johnson (2013: Chapter 6) call “cost sensitive training.” Related thinking can be found in focal loss approaches (Lin et al., 2017) and implicitly in gradient boosting (Friedman, 2001). Rednick and his colleagues (2021) offer the most involved approach that weights more heavily the hard to classify observations, and a specified outcome class can be given a higher priority as well.222222 We coded their loss function in keras3 in R. It ran, but did far worse that the usual cross-entropy loss, perhaps because the classifier was already too successful with the within-subject design and the balanced data that resulted. There were too few false negatives and false positives usefully to weight. We also found tuning a conceptual challenge when we tried to return to a highly imbalanced source population response variable combined with a plausible stakeholder consensus cost ratio. In their classic book, Luce and Raiffa (1957) formulate the problem as risk minimization rather than loss minimization. Broadly framed, asymmetric loss functions are required to account for asymmetric fitting costs.
@@
The conflation of corrections for sample imbalance with the estimated misclassification cost ratio is easily illustrated. Using the expressions above for the weights, the PNW classification exercise was repeated. Recall that the response variable was binary, and there were the same number of grid cells that experienced true heat wave precursors as experienced faux heat wave precursors. Using the month of June, 2021 as the target population, for a little more than 13% of the days, the 144 grid cells experienced the extreme heat wave precursors. For the other days, the 144 grid cells the precursors were for a faux heat wave. In the data analyzed, for 50% of the days, the 144 grid cells experienced extreme heat wave precursors. A disparity between 13% and 50% may be more extreme than anticipated by Manski and Lerman (1977), but we can proceed for illustrative purposes.
@@
After weighting, the population imbalance was approximately reproduced. Before weighting, the classification error rate was a little over 5%, and there were nearly the same number of false negatives as false positives. The cost ratio was about 1.0, conveying equal costs. When Manski-Lerman weighing was applied, the classification error rate rose to nearly 14%, and the cost ratio of false positives to false negatives increased from about 1.0 to nearly 10.0. False positives were now almost 10 times more costly than false negatives. In other words, the costs of heat wave false alarms were substantially up-weighted compared to the costs of true alarms not sounded. Many stakeholders might strongly prefer the reverse. In short, weighting changes the prior distribution of the response variable and the relative costs of false negatives to false positives.
@@
For the classification results reported for the PNW and the greater Phoenix area, we had no access to stakeholders nor any concrete policy context. The numbers of false negatives and false positives were very few and of roughly equal size. Classifier training determined an empirical cost ratio of about 1 to 1. Going forward in a real policy setting combined with stakeholder input, the within-subject research design could be revised to better capture practical concerns. The main requirement is that grid cells continue to capitalize on the within-subject structure. Additional faux heat waves could easily be added so that a more appropriate, but manageable, imbalance was achieved. Another possibility might be to train the classifier on balanced data but to improve fitted values, use as input an altered dataset with many more faux heat waves.